OpenStudy (anonymous):

An architect designs a theater with 18 seats in the first row, 22 in the second, 26 in the third, and so on. If the theater is to have a seating capacity of 1120, how many rows must the architect use in his design?

OpenStudy (anonymous):

18+x(18+4)=1120 x=50 , i think it is 51 row

OpenStudy (anonymous):

its not, unfortunately. I have online homework and it says its wrong. But thanks for trying!

OpenStudy (anonymous):

18+n4+18n=1120 22n=1120 n=560

OpenStudy (anonymous):

561 row i thinkl

OpenStudy (anonymous):

oh wrong

hero (hero):

n = 20

hero (hero):

I was way off the first time

OpenStudy (anonymous):

i will try to use pyton

OpenStudy (aravindg):

n=20

OpenStudy (aravindg):

i got it

OpenStudy (aravindg):

32x+4x^2-2240=0

OpenStudy (anonymous):

yes 20

OpenStudy (anonymous):

yes 20 here the python code: n=0 Y=0 y=0 while 18+n*18+y<1120: print(18+n*18+y) n+=1 y=Y+4*n Y=y else: print(n) print(y) print(18+n*18+y)

OpenStudy (anonymous):